Solve the following pairs of simultaneous equations.
x/2 + y/3 = 5
x - y = 10

Answer by checkley79(3341)

x/2 + y/3 = 5 multiply by 3 & add.
x - y = 10
3x/2+y=15 add.
-------------
5x/2=25
5x=25*2
5x=50
x=50/5
x=10 ans.
10-y=10
-y=10-10
-y=0
y=0 ans.
Proof:
10/2+0/3=5
5=5